ergy & Work A 1.12-kg SkyBall is thrown straight down with a speed of 16.81 m/s from a height of 1.0 m. If no energy is lost in the bounce, how high up does the ball go? Answer: check Incorrect Marks for this submission: 0.00/1.00 us pageFinish attempt
Answer #1

here,


the mass of ball ,m = 1.12 kg

initial speed , u = 16.81 m/s

the initial height , h0 = 1 m

let the maximum height gained be h

using conservation of energy

0.5 * m * u^2 + m * g * h0 = m * g * h

0.5 * 15.81^2 + 9.81 * 1 = 9.81 * h

solving for h

h = 13.74 m

the maximum height gained is 13.74 m